OpenStudy (kpop4life123):

What is the product of 3.4 × 10–14 and 1.8 × 1028? (5 points) 6.12 × 1042 5.2 × 1042 6.12 × 1014 9.4 × 1014

OpenStudy (cggurumanjunath):

C Option

OpenStudy (unklerhaukus):

\[(3.4 \times 10^{–14}) \times (1.8 × 10^{28})\\ =(3.4\times1.8)\times10^{-14+28}\\ =\]